# A. CamillettiAI weather from Euro-Atlantic regimes

### v1

@@ Appendix A Four winter and summer WR
We compute the four WR for winter and summer using the same methodology described in Sec. 3.1, only changing the number of centroids used to cluster the data. The mean $Z500$ field (black isocontours) and the corresponding average $Z500$ anomalies during winter (December, January, February) and summer (June, July, August), associated with the four Euro-Atlantic WR, are presented in Fig. A.1 and Fig. A.2, respectively. The four WR are the ZO, ScBL, AT and GL. The two WR resemble the two phases of the NAO, i.e. the ZO and the GL.
@@
Figure A.1: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in winter (December, January, February), associated with the four Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The ”No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [Grams et al., 2017].
@@
Figure A.2: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in summer (June, July, August), associated with the four Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The ”No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [Grams et al., 2017].
Appendix B Differences in the WR computations with respect to Grams et al. [2017]
@@
We now highlight the major differences to the procedure described in Grams et al. [2017] and in Büeler et al. [2021]. The authors of the mentioned work used the ERA-Interim reanalysis [Dee et al., 2011] six-hourly 500 hPa GPH from 11 January 1971 to 31 December 2015; we instead use ERA5 daily 500 hPa GPH from 1st of January 1940 to 31 December 2024 . Moreover, the anomalies computation, the EOF analysis, the K-means clustering, and the definition of the cluster’s mean field are performed only in the climatology period. Grams et al. [2017] applied 90- and 30-day moving window averages to smooth the calendar day mean and standard deviation. In contrast, we use a 15-day centered moving average for each calendar day. We observed that using window sizes larger than approximately 30 days tends to preserve seasonal trends in the resulting anomalies. This explains the minor difference that can be found when comparing Fig. 1 with the corresponding Supplementary Figure 1 in Grams et al. [2017] and Figure 1 in Büeler et al. [2021].
Appendix C Anomaly reconstruction without indices
@@
Figure C.1: Spatial ACC (top) and CE (bottom) of the AI model reconstruction of the winter mean two-meter temperature and total precipitation anomalies without using the information encoded in the WR indices. The skills are computed in the test years (from January 2011 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC higher than 0.5.
@@
As a baseline for evaluating the added value of the WR indices, we trained the same AI model described in Sec. 4 to reconstruct monthly mean two-meter temperature and total precipitation anomalies without including the WR indices as input. The obtained spatial ACC and CE for the winter mean two-meter temperature and total precipitation anomalies are shown in Fig. C.1. Despite the absence of synoptic-scale information, the model exhibits low but positive ACC in certain European regions, likely due to its ability to learn and exploit the underlying climate trend present in the training data. Nevertheless, the CE value is around zero or negative, indicating that the model is not capable of replicating the amplitude of the fluctuations in the test dataset.
Appendix D Reconstruction excluding the validation dataset
In the main text, we presented the reconstruction skills the AI-model for the summers and winters from 2011 to 2024 , including the validation dataset form January 2011 to December 2018. The inclusion of the validation dataset is necessary to ensure a statistically relevant evaluation, as the independent test set alone is too limited in duration to provide robust skill estimates. The validation dataset is used to monitor the training and to stop it before overfitting occur, in order to guarantee optimal generalization of the AI-model. Therefore, including the validation dataset onto the test period can introduce a bias in the results, possibly improving the model’s performance. To assess the magnitude of this effect, we re-evaluate the reconstruction performance using only the independent test period, from January 2019 to 2024 .
Figure D.2: AI model spatial MAE (top), ACC (middle) and CE (bottom) in reconstructing the winters (first and third columns) and summers (second and fourth columns) mean two-meter temperature and total precipitation anomalies using seven WR indices in the test years (from January 2019 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC (CE) higher than 0.5 (0.3).
Fig. D.2 shows the spatial distribution of the MAE, ACC and CE for the monthly mean two-meter temperature and total precipitation anomaly in 2019 – 2024, reconstructed using the seven year-round WR indices. This figure is the analogue of Fig. 3 and exhibits similar spatial patterns. As described in Sec. 5.1, model skill is consistently higher in winter and in central/western Europe, with temperature reconstruction being better than precipitation. The spatial patterns of high (low) ACC and CE (MAE) are largely shared between the two figures, though some quantitative differences are evident. In particular, where higher (lower) ACC and CE values are found, the AI-model tested on the shorter dataset scores better (worse) than the same AI-model tested on the extended dataset, especially for the two-meter temperature. The shorter validation period also limit the statistical significance of the correlations, as demonstrated by the large areas with p-value lower than 0.05 (less than 95% confidence). The p-value test is less reliable in discriminating significant from nonsignificant correlations in the regions with high or low ACC. or CE, which is computed as a temporal mean over the evaluation period (Eq. 5), using a shorter time window can exaggerate extreme values. This leads to larger differences in CE compared to the full-period evaluation, as reflected in Fig. 3.
two-meter temperature total precipitation
MAE ACC CE MAE ACC CE
winter 0.02 0.02 0.02 -0.01 -0.01 0.03
summer -0.04 0.04 0.16 -0.01 0.03 0.07
Table 2: Difference between medians of MAE, ACC and CE of the two-meter temperature and total precipitation reconstructions evaluated including the validation dataset (2011 - 2011 ) and excluding it (2019 - 2024 ). The skills are evaluated separately for winter (DJF) and summer (JJA). The model use the seven WR indices as input.
Tab. 2 presents the differences between the median skill metrics calculated with and without the inclusion of the validation dataset. The difference are typically small, of the order of few percent, indicating that including the validation data do not substantially bias in the results. The only notable increase occurs in the CE for summer temperature, likely because it is not bounded from below and degrades heavily when computed over fewer years. Excluding the validation data leads to more extreme negative CE values in certain regions, further amplifying discrepancies relative to the full dataset.
In conclusion, a comparison of model performance over the full evaluation period (2011–2024), which includes the validation years, and the independent test period (2019–2024) reveals reveals no consistent signs of overfitting or significant bias due to inclusion of the validation dataset.

### v2

@@ Appendix A Four winter and summer WR
We compute the four WR for winter and summer using the same methodology described in Sec. 3.1, with the only change being the number of centroids used to cluster the data. The mean $Z500$ field (black isocontours) and the corresponding average $Z500$ anomalies during winter (December, January, February) and summer (June, July, August), associated with the four Euro-Atlantic WR, are presented in Fig. A.1 and Fig. A.2, respectively. The four WR are the ZO, ScBL, AT and GL.
@@
Figure A.1: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in winter (December, January, February), associated with the four Euro-Atlantic WR. The average is performed in the 1981-2010 climatology.
@@
Figure A.2: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in summer (June, July, August), associated with the four Euro-Atlantic WR. The average is performed in the 1981-2010 climatology.
Appendix B Differences in the WR computations relative to Grams et al. (2017)
@@
We now highlight the major differences to the procedure described in grams_balancing_2017 and in bueler_year-round_2021. The authors of the mentioned work used the ERA-Interim reanalysis [dee_era-interim_2011] six-hourly 500 hPa GPH from 11 January 1971 to 31 December 2015; we instead use ERA5 daily 500 hPa GPH from 1st of January 1940 to 31 December 2024 . Moreover, the anomalies computation, the EOF analysis, the K-means clustering, and the definition of the cluster’s mean field are performed only in the climatology period. [grams_balancing_2017] applied 90- and 30-day moving window averages to smooth the calendar day mean and standard deviation. In contrast, we use a 15-day centered moving average for each calendar day. We observed that using window sizes larger than approximately 30 days tends to preserve seasonal trends in the resulting anomalies. This explains the minor difference that can be found when comparing Fig. 1 with the corresponding Supplementary Figure 1 in grams_balancing_2017 and Figure 1 in bueler_year-round_2021.
Appendix C Training setup and hyperparameters
The models are trained using mean squared error (MSE) as the loss function (see Sec. 4.2). Training is performed using the Adam optimizer [kingma2017adammethodstochasticoptimization] with mini-batches and early stopping based on the minimum validation loss, so the total number of epochs varies between runs and is not fixed a priori. We employ a learning-rate plateau scheduler: whenever the validation loss does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 0.5. The network includes dropout with a dropout rate of 0.2 to mitigate overfitting. Other training hyperparameters are reported in Tab. C.0 Within reasonable ranges, hyperparameters such as batch size, initial learning rate, patience, and dropout rate have a limited impact on the final performance of the models. In our experiments, variations in these quantities result in changes comparable to the run-to-run variability due to randomness (see Tables 1 and 2).
Table C.0: Summary of training hyperparameters used in this work.
Hyperparameter Value
Loss function MSE
Optimizer Adam
Initial learning rate $10^{-3}$
LR reduction factor 0.5
LR plateau patience 10 epochs
Batch size 8
Dropout rate 0.2
Appendix D Anomaly reconstruction without indices
@@
Figure D.1: Spatial ACC (top) and CE (bottom) of the AI model reconstruction of the winter mean two-meter temperature and total precipitation anomalies without using the information encoded in the WR indices. The skills are computed in the test years (from January 2011 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC higher than 0.5.
@@
As a baseline for evaluating the added value of the WR indices, we trained the same AI model described in Sec. 4 to reconstruct monthly mean two-meter temperature and total precipitation anomalies without including the WR indices as input. The obtained spatial ACC and CE for the winter mean two-meter temperature and total precipitation anomalies are shown in Fig. D.1. Despite the absence of synoptic-scale information, the model exhibits low but positive ACC in certain European regions, likely due to its ability to learn and exploit the underlying climate trend present in the training data. Nevertheless, the CE value is close to zero or negative, indicating that the model is unable to replicate the amplitude of the fluctuations in the test dataset.
Appendix E Evaluation of the SEAS5 WR forecast
Figure E.1: CRPS (lower is better) and ensemble mean correlation (higher is better) of SEAS5 WR indices forecast, $I_{\rm wr}^{\rm SEAS5}$, with respect to ERA5 WR indices, $I_{\rm wr}$. The labels refer to the WR names: AT, ZO, ST, AR, EuBL, ScBL, GL.
As noted in Sec. 5.5, the low skills of the AI model driven by the SEAS5 forecasted WR indices, $I_{\rm wr}^{\rm SEAS5}$, suggest that the SEAS5 forecast of the WR indices is not very skillful. Fig. E.1 show the CRPS (lower is better) and the ensemble mean correlation (higher is better) between $I_{\rm wr}^{\rm SEAS5}$ and $I_{\rm wr}$, i.e. between the monthly indices forecasted by the bias corrected SEAS5 and the indices computed from ERA5 Z500 (see Sec. 3.1 and Sec. 3.4). Both the CRPS and the ensemble mean correlation are good in the initialization month, but the skill degrades abruptly from the second forecasted month. As usually done in the literature, we initialize SEAS5 in November for winter forecasts (DJF), and in May for summer forecasts (JJA). The considered forecast is relative to the second, third, and fourth forecasted months, which, in general, has poor skills.